OpenStudy (anonymous):

The weight of 10,000 identical samples of a substance is 10 pounds. What is the weight of 10 samples?

OpenStudy (akashdeepdeb):

The price of 10 candies is 1 pound. What is the price of one candy?

OpenStudy (akashdeepdeb):

@ag1234

OpenStudy (anonymous):

idk

OpenStudy (akashdeepdeb):

The price of 10 candies is 100 pounds. What is the price of one candy?

OpenStudy (akashdeepdeb):

@ag1234 ?

OpenStudy (anonymous):

$10???

OpenStudy (akashdeepdeb):

correct.

OpenStudy (anonymous):

yayay!!!

OpenStudy (akashdeepdeb):

Now how did you calculate that?

OpenStudy (anonymous):

100 divided by 10???

OpenStudy (akashdeepdeb):

Very good. Now what if the price of those 10 candies were 1 dollar. What do you think would be the price of 1 candy?

OpenStudy (anonymous):

$1???????????

OpenStudy (akashdeepdeb):

nO. See in the previous case, it was 10 candies = 100 dollars and you found out the price of one candy by 1 candy = 100/10 So when 10 candies = 1 dollar then the price of one = ?? One candy = 1/10 dollars!! :D

OpenStudy (akashdeepdeb):

Do you know fractions or decimals?

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

Did you CLEARLY understand that up there? ^

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait how does this help me with what im doing?

OpenStudy (akashdeepdeb):

No probs I'll explain again. You have a similar kind of problem to solve. :)

OpenStudy (anonymous):

okayy

OpenStudy (akashdeepdeb):

Let us take it like this: Let 10 books be 10 dollars.